OpenStudy (anonymous):

√x-2 = x-8 my options are {11}, {6,11},{}

OpenStudy (anonymous):

what is the question? i'm confused

OpenStudy (anonymous):

The \[\sqrt{x-2} = x-8\]

OpenStudy (anonymous):

By squaring right side: (x-2) = (x-8)^2 (x-2) = (x)^2 - 2(x)(8) + (-8)^2 (x-2) = x^2 - 16x + 64 0 = x^2 -17x + 66 (x-11)(x-6) = 0 x=? and ?

OpenStudy (anonymous):

it's the {6,11}